If
α=3i−j+2k,

β=2i+j−k and
,
γÌ…=i−2j+2k, find

α×β×γ.
1
Expert's answer
2020-06-15T19:11:51-0400

we know that if

"\overrightarrow{\alpha}" "= x_1\hat{i}+y_1\hat{j}+z_1\hat{k}" and "\overrightarrow{\beta}=x_2\hat{i}+y_2\hat{j}+z_2\hat{k}"

then "\overrightarrow{\alpha }" x "\overrightarrow{\beta} = \begin{vmatrix}
 \hat{i} & \hat{j} &\hat{k} \\
 x_1 & y_1 &z_1\\
x_2&y_2&z_2
\end{vmatrix}"


Here "\overrightarrow{\alpha} = 3\hat{i}-\hat{j}+2\hat{k}" , "\overrightarrow{\beta}=2\hat{i}+\hat{j}-\hat{k}" and "\overrightarrow{\gamma}= \hat{i}-2\hat{j}+2\hat{k}"

so,

"\overrightarrow{\alpha}" x "\overrightarrow{\beta} = \begin{vmatrix}
 \hat{i} & \hat{j} &\hat{k} \\
 3 & -1 &2\\
2&1&-1
\end{vmatrix}"

expanding along Row 1.


"\overrightarrow{\alpha}" x "\overrightarrow{\beta}= \hat{i}\begin{vmatrix}
 -1 & 2 \\
 1 & -1
\end{vmatrix}-\hat{j}\begin{vmatrix}
 3 & 2 \\
 2 & -1
\end{vmatrix}+\hat{k}\begin{vmatrix}
 3 & -1 \\
 2 & 1
\end{vmatrix}"


"= \hat{i}(1-2)-\hat{j}(-3-4)+\hat{k}(3+2)"

"=-\hat{i}+7\hat{j}+5\hat{k}" .


Following the same way "\overrightarrow{\alpha}" x "\overrightarrow{\beta}" x "\overrightarrow{\gamma}" can be determined .


so ,

"\overrightarrow{\alpha}" x "\overrightarrow{\beta}" x "\overrightarrow{\gamma}=" "\begin{vmatrix}
 \hat{i} & \hat{j} &\hat{k} \\
 -1 & 7 &5\\
1&-2&2
\end{vmatrix}"

"=\hat{i}\begin{vmatrix}
 7& 5 \\
 -2 & 2
\end{vmatrix}-\hat{j}\begin{vmatrix}
 -1 & 5 \\
 1 & 2
\end{vmatrix}+\hat{k}\begin{vmatrix}
 -1 & 7 \\
 1 & -2
\end{vmatrix}"

"=\hat{i}(14+10)-\hat{j}(-2-5)+\hat{k}(2-7)"

"=24\hat{i}+7\hat{j}-5\hat{k}"


The answer is :

"\therefore \overrightarrow{\alpha}" x "\overrightarrow{\beta}" x "\overrightarrow{\gamma}=24\hat{i}+7\hat{j}-5\hat{k}."